OpenStudy (anonymous):

Complete the table by finding the missing polynomial. Enter the polynomial in standard form. Polynomial 1 --Polynomial 2--- Sum 17x + 9 ? 20x +17 AP 1

zepdrix (zepdrix):

Hey Billy! We have a Sum. Polynomial 1 + some other polynomial gives us the result in the Sum column. \[\large\rm (17x+9)+(?x+?)=20x+17\]Simply equate like-terms,\[\large\rm 17x+?x=20x\]How many x's are we missing?\[\large\rm 9+?=17\]And how many of these units are we missing?

OpenStudy (faiqraees):

lets suppose your second polynomial is\[\large\rm ax+b\] so adding the first polynomial with the second polynomial would give an equation like this \[\large\rm ax+b+17x+9=20x+17\]\[\large\rm ax+b=20x+17x-17-9\]\[\large\rm ax+b= 37x-26\] there you have the second polynomial.